A new water pump has a capacity of 60 m3/day. If its capacity goes down by 15% every year, in how many years will the capacity go down to 20 m3/day? -Differential Equation Problem
Answer by Alan3354(69443)

A new water pump has a capacity of 60 m3/day. If its capacity goes down by 15% every year, in how many years will the capacity go down to 20 m3/day?
--------------
Each year its capacity is 85% of the previous year.
C = 60*0.85^t --- t in years
20 = 60*0.85^t
1/3 = 0.85^t
log(1/3) = t*log(0.85)
t = log(1/3)/log(0.85)
t =~ 6.76 years